OpenStudy (anonymous):

what is the integration of 4x^2-2x+2/4x^2-4x+3

OpenStudy (anonymous):

\[\int\limits (4x^2-2x+2/4x^2-4x+3) dx\]

OpenStudy (dumbcow):

first step would be use long division

OpenStudy (dumbcow):

\[\rightarrow \int\limits_{}^{}1 + \frac{2x-1}{4x^{2}-4x+3} dx\] then make substitution u = 4x^2 -4x +3 du = 8x-4 = 4(2x-1) dx \[\rightarrow x +\frac{1}{4}\int\limits_{}^{}\frac{1}{u} du\]

OpenStudy (anonymous):

thanks how did you do long division?

OpenStudy (dumbcow):

1 + (2x-1)/(4x^2-4x+3) ___________________ 4x^2 -4x +3 | 4x^2 -2x +2 -(4x^2 -4x +3) -------------- 2x -1

OpenStudy (anonymous):

oh k thats how it works